OpenStudy (anonymous):

How are you supposed to reason when solving the equation tanx=1/sqrt3?

OpenStudy (jack1):



OpenStudy (anonymous):

Yes, but what does this say about x?

OpenStudy (anonymous):

I guess that it would say that the hypotenuse is 2, so therefore you can use sin or cos in order to calculate the x?

OpenStudy (anonymous):

Meaning what is the solve for the system {sin(x)=1/2;cos(x)=sqrt3/2}

OpenStudy (anonymous):

Which would conclude the solution x=Pi/6+2nPi, for any integer n.

OpenStudy (anonymous):

Do you need to reason this way? Is there a simpler way that always yields the correct solution like this way?

OpenStudy (anonymous):

OOOOOH, I see where I went wrong. :D

OpenStudy (anonymous):

I expressed sin^2x in terms of tan^2x; this means that |sinx|=1/2 and therefore gave a false solution.

OpenStudy (anonymous):

Thanks for your simple yet mind-opening triangle though, Jack! :D

OpenStudy (unklerhaukus):